Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 86     Entry time: 17 Mar 2023, 23:43
Author: Robert, Zhong, Manuel 
Group: SCS 
Subject: Late shift summary 17.03.2023 

Activities:

  • continued tr-RIXS scans
  • tried to find ways to mitigate the energy drifts of the mono (see below)
  • confirmed in user data that the FEL-PPL timing was not changing (see below)
  • setting up for tr-RIXS at beginning of night shift

Problems:

  • the mono energy has been drifting, presumably due to warming of the grating.
  • automatic correction of the elastic line in the RIXS data was not working reliably
  • we suspected a timing drift, but it seems that the monitor (photodiode + oscilloscope) itself is drifting. That's a new one.
ELOG V3.1.4-7c3fd00